OpenStudy (anonymous):

Help with this parabola, 3x^2 +4x+12

OpenStudy (anonymous):

@Nnesha

OpenStudy (campbell_st):

well I can tell you its positive definite

Nnesha (nnesha):

statement ? factor ? solve for x ? graph ? write in vertex form? find vertex point ? what is axis sym ?

OpenStudy (anonymous):

I dont know how to find the zeros

Nnesha (nnesha):

alright zeros(solutions , x-intercept) you can use quadratic formula `OR` you can factor it

Nnesha (nnesha):

which one is easy for you ? factors or quadratic formula ?

OpenStudy (anonymous):

Factors

Nnesha (nnesha):

ohh well nvm you have to apply the quadratic formula \[\huge\rm x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

Nnesha (nnesha):

ax^2+bx+c quadratic equation where a=leading coefficient b=middle term c-constant term so what is a b and c in that equation ?

Nnesha (nnesha):

\(\huge\color{reD}{\rm b^2-4ac}\) `Discriminant` you can use this to find if the equation is factorable or not if ` b^2-4ac > 0` then there are 2 real zeros if ` b^2-4ac = 0` then there is one real root if ` b^2-4ac < 0` then you will get two complex roots (no -x-intercept)

Nnesha (nnesha):

zeros is another word they use for solutions and x-intercept

OpenStudy (lynfran):

can u show me how to write it in vertex form .. @Nnesha

Nnesha (nnesha):

complete the square y=a(x-h)^2+k ?

OpenStudy (lynfran):

o thats all lol... im now doing this ..thanks

Nnesha (nnesha):

i mean convert 3x^2 +4x+12 into y=a(x-h)^2+k <---(vertex form) b using completing the square method :=)

OpenStudy (lynfran):

i know what u meant thanks

Nnesha (nnesha):

np :)

OpenStudy (lynfran):

so it would be 3x^2+(4x/2)^2+......+12-..... (3x^2+4x+4)+12-4 (3x+2)^2+8 is this correct

OpenStudy (lynfran):

i think i did something wrong

Nnesha (nnesha):

okay you need to complete the square of (3x^2+4x) it would be great if you take out the 3

Nnesha (nnesha):

i've to go rn cya later

OpenStudy (lynfran):

ok